OpenStudy (anonymous):

Jane is organizing a fundraiser to buy a ping-pong table for the community center. The table costs $500.00. Jane is asking contributors to pay for an equal share of the cost of the table. She already has five contributors lined up. What function models the cost per share, C, with respect to the number of contributors, n? How many contributors must Jane recruit in order for the cost of each share to be $20.00?

OpenStudy (smokeysthename):

@hartnn can you help us with this question?

hartnn (hartnn):

one table = 500$ one contributor's share =C so., 5 contributor's share =5C n contributor's share =nC

OpenStudy (anonymous):

Is that the answer? or are you explaining it step by step ?

OpenStudy (smokeysthename):

i think he's explaining it step by step

OpenStudy (anonymous):

Okay cool.

hartnn (hartnn):

so, we have nC=500 which gives us , cost per share C =500/n.

OpenStudy (anonymous):

Okay @hartnn i see where your going. Is that my answer ?

hartnn (hartnn):

for 2nd part, you want C= 20, so 20=500/n find n

hartnn (hartnn):

yes, that was 1st part answer.

OpenStudy (smokeysthename):

okay so would you multiply 20 and 500 ?

OpenStudy (smokeysthename):

or would it be 500 divided by 20

OpenStudy (anonymous):

Yeah and 500/20 will equal 25

hartnn (hartnn):

20=500/n n=500/20 = 25

hartnn (hartnn):

yes, she already had 5 contributors, so she will need 20 more.

OpenStudy (smokeysthename):

so thats the answer?

OpenStudy (anonymous):

Thats what im thinking

OpenStudy (phi):

Smokey, if 25 people pay $20 how much money do they give?

OpenStudy (smokeysthename):

500

OpenStudy (anonymous):

Thats what i got because i did 25*20 to queal 500

hartnn (hartnn):

so, you got 500, right ? yes, that was your answer.

OpenStudy (anonymous):

Great :D

OpenStudy (anonymous):

Snip answer @SmokeysTheName

OpenStudy (smokeysthename):

k one sec

OpenStudy (anonymous):

Got cha

hartnn (hartnn):

ask if there are any doubts...

OpenStudy (anonymous):

Okay @hartnn Thanks

OpenStudy (smokeysthename):

thank you @hartnn

OpenStudy (smokeysthename):

@Kayy_Drizzyy

hartnn (hartnn):

welcome to both :)

OpenStudy (smokeysthename):

can you answer another one @hartnn

hartnn (hartnn):

i'll try.

OpenStudy (smokeysthename):

ok ill tag you in it.

OpenStudy (anonymous):

I got the same answer

OpenStudy (anonymous):

did you push finish @SmokeysTheName ?

OpenStudy (smokeysthename):

nooo we still havent figured out the hair dye one w/ percentages

OpenStudy (anonymous):

Ohhhhhh yeah lets ask @hartnn for help on that one.

OpenStudy (smokeysthename):

im getting help on that one before i submit it.

OpenStudy (smokeysthename):

i did ill tag you in my question

OpenStudy (anonymous):

great :)